# Regression-Based Single-Point Zeroth-Order Optimization

### v1

@@ A.2 Proofs for Lemma 1, Theorem 1, and Theorem 2
Suppose Assumption 1 holds. Suppose $2\leq m=\mathcal{O}(d)$. Consider any $t\geq m$. Suppose that
@@
Then, by picking $\eta=\mathcal{O}\left(\frac{1}{dL}\right)$, we can ensure that
@@
Next, we restate and prove Lemma 1, which shows that for any $10<c<$ 1, we can inductively guarantee that the gradient estimation bias $\|\xi_{t}\|$ always satisfies $\|\xi_{t}\|\leq c\|\nabla f(\bm{x}_{t})\|$ for all $t\geq m$.
@@
holds for all $j<t$. Then, for any $10<c<$ 1, by picking $\eta=C\frac{1}{dC_{d}L}$ for some sufficiently small absolute constant $C>0$, we will show that the inductive claim implies that
@@
=$\stackrel{{\scriptstyle\textnormal{(vi)}}}{{\mathstrut{=}}}$ $\ \frac{C_{d}L}{2}\sum_{j=1}^{m}4\|\bm{x}_{t-j+1}-\bm{x}_{t-j}\|$
@@
where (A.2) holds by $L$-smoothness of $f$, (A.2) holds by definition of $\hat{\bm{x}}_{t}:=\hat{\bm{x}}_{t}+\eta_{t-1}\|\bm{g}_{t-1}\|\bm{u}_{t}$ due to the adaptive selection of the smoothing radius $\delta_{t}:=\eta\|\bm{g}_{t-1}\|$ and the fact that each $\bm{u}_{t}$ has unit norm, (A.2) utilizes Assumption 2, (A.2) again uses the fact that each $\bm{u}_{t}$ has a unit norm, (A.2) holds by our assumption at the start of this subsection that $C_{d}\geq 1$, (A.2) holds since $\bm{x}_{t-j+1}=\bm{x}_{t-j}+\eta_{t-j}\bm{g}_{t-j}$, and (A.2) follows by our inductive assumption.
@@
Now, (recalling that $m=\mathcal{O}(d)$), by picking $\eta=\frac{C}{dC_{d}L}$ for a sufficiently small $C>0$, we may conclude that
@@
Summing this from $t=m$ to $t=T$ and plugging in our choice of $\eta=\mathcal{O}\left(\frac{1}{dC_{d}L}\right)$, we observe that
$\ f(\bm{x}_{T})-f(\bm{x}_{m})\leq-\mathcal{O}\left(\frac{1}{dC_{d}L}\sum_{t=m}^{T-1}\|\nabla f(\bm{x}_{t})\|^{2}\right)$
$\implies$ $\ \mathcal{O}\left(\frac{1}{dC_{d}L}\sum_{t=m}^{T-1}\|\nabla f(\bm{x}_{t})\|^{2}\right)\leq f(\bm{x}_{m})-f(\bm{x}_{T})$
$\implies$ $\ \frac{1}{T}\sum_{t=m}^{T-1}\|\nabla f(\bm{x}_{t})\|^{2}\leq\mathcal{O}\left(\frac{dC_{d}L\left(f(\bm{x}_{m})-f^{*}\right)}{T}\right),$
@@
Appendix B Empirical Study of $C_{d}$ in Assumption 2
@@ B.1 Ridge Regression
Figure 4: Convergence of L-RESZO on the ridge-regression problem in (26) for dimensions $d=100,400,900$. The $C_{d}$ ratios shown in Figure 3 are based on the algorithmic trajectories here.
B.2 Logistic Regression

### v2

@@ B.2 Proofs for Lemma 1, Theorem 1, and Theorem 2
Suppose Assumption 1 holds. Suppose $m>d$. Consider any $t\geq m$. Suppose that
@@
Then, by picking $\eta=\mathcal{O}\left(\frac{1}{mL}\right)$, we can ensure that
@@
Next, we restate and prove Lemma 1, which shows that for any $0<c<1$, we can inductively guarantee that the gradient estimation bias $\|\xi_{t}\|$ always satisfies $\|\xi_{t}\|\leq c\|\nabla f(\bm{x}_{t})\|$ for all $t\geq m$.
@@
holds for all $j<t$. Then, for any $0<c<1$, by picking $\eta=C\frac{1}{mC_{d}L}$ for some sufficiently small absolute constant $C>0$, we will show that the inductive claim implies that
@@
$\stackrel{{\scriptstyle\textnormal{(vi)}}}{{\mathstrut{=}}}$ $\ \frac{C_{d}L}{2}\sum_{j=1}^{m}4\|\bm{x}_{t-j+1}-\bm{x}_{t-j}\|$
@@
where (B.2) holds by $L$-smoothness of $f$, (B.2) holds by definition of $\hat{\bm{x}}_{t}:=\hat{\bm{x}}_{t}+\eta_{t-1}\|\bm{g}_{t-1}\|\bm{u}_{t}$ due to the adaptive selection of the smoothing radius $\delta_{t}:=\eta\|\bm{g}_{t-1}\|$ and the fact that each $\bm{u}_{t}$ has unit norm, (B.2) utilizes Assumption 2, (B.2) again uses the fact that each $\bm{u}_{t}$ has a unit norm, (B.2) holds by our assumption at the start of this subsection that $C_{d}\geq 1$, (B.2) holds since $\bm{x}_{t-j+1}=\bm{x}_{t-j}+\eta_{t-j}\bm{g}_{t-j}$, and (B.2) follows by our inductive assumption.
@@
Now, by picking $\eta=\frac{C}{mC_{d}L}$ for a sufficiently small $C>0$, we may conclude that
@@
We next restate and prove Lemma 2.
See 2
Proof.
Consider any positive integer $t\geq m$. Fix any $0<c<1$. By our assumptions, applying Lemma 1, we find that there exists a sufficiently small absolute constant $C>0$ such that when $\eta=\frac{C}{mC_{d}L}$, we have
$\|\xi_{t}\|\leq c\|\nabla f(\bm{x}_{t})\|$
for any $t\geq m$.
Next, observe that
$\ f(\bm{x}_{t+1})-f(\bm{x}_{t})$
$=$ $\ -\eta\langle\nabla f(\bm{x}_{t}),\nabla f(\bm{x}_{t})\rangle-\eta\langle\nabla f(\bm{x}_{t}),\bm{g}_{t}-\nabla f(\bm{x}_{t})\rangle$
$\quad\quad+\frac{L}{2}\|\eta(\bm{g}_{t}-\nabla f(\bm{x}_{t})+\nabla f(\bm{x}_{t}))\|^{2}$
$\stackrel{{\scriptstyle\textnormal{(ii)}}}{{\mathstrut{\leq}}}$ $\ -\eta\|\nabla f(\bm{x}_{t})\|^{2}+\frac{\eta}{2}\left(\|\nabla f(\bm{x}_{t})\|^{2}+\|\xi_{t}\|^{2}\right)$
$\quad\quad+\eta^{2}L\left(\|\nabla f(\bm{x}_{t})\|^{2}+\|\xi_{t}\|^{2}\right)$
$\leq$ $\ -\frac{\eta}{4}\|\nabla f(\bm{x}_{t})\|^{2}+\eta\|\xi_{t}\|^{2},$ (36)
where (B.2) follows by $L$-smoothness of $f$, (B.2) follows by application of Young’s equality as well as the notation $\xi_{t}:=\bm{g}_{t}-\nabla f(\bm{x}_{t})$, and the final inequality follows by choosing $\eta\leq\frac{1}{4L}$.
$\|\xi_{t}\|\leq\frac{1}{\sqrt{8}}\|\nabla f(\bm{x}_{t})\|,$
we find that
$\leq$ $\ -\frac{\eta}{8}\|\nabla f(\bm{x}_{t})\|^{2},$
as desired.
@@
Summing this from $t=m$ to $t=T$ and plugging in our choice of $\eta=\mathcal{O}\left(\frac{1}{mC_{d}L}\right)$, we observe that
$\ f(\bm{x}_{T})-f(\bm{x}_{m})\leq-\mathcal{O}\left(\frac{1}{mC_{d}L}\sum_{t=m}^{T-1}\|\nabla f(\bm{x}_{t})\|^{2}\right)$
$\implies$ $\ \mathcal{O}\left(\frac{1}{mC_{d}L}\sum_{t=m}^{T-1}\|\nabla f(\bm{x}_{t})\|^{2}\right)\leq f(\bm{x}_{m})-f(\bm{x}_{T})$
$\implies$ $\ \frac{1}{T}\sum_{t=m}^{T-1}\|\nabla f(\bm{x}_{t})\|^{2}\leq\mathcal{O}\left(\frac{mC_{d}L\left(f(\bm{x}_{m})-f^{*}\right)}{T}\right),$
@@
Appendix C Empirical Study of $C_{d}$ in Assumption 2
We assume throughout this section that $m=\mathcal{O}(d)$, so we may denote $C_{d}:=C_{d,m}$.
@@ C.1 Ridge Regression
Figure 4: Convergence of L-ReSZO on the ridge‐regression problem in (27) for dimensions $d=100,400,900$. The $C_{d}$ ratios shown in Figure 3 are based on the algorithmic trajectories here.
C.2 Logistic Regression